find two supplementary angles such that the measure of the first angle is 60 degrees less than twice the second?

Answer by checkley77(12569)

x+y=180 or x=180-y
x=2y-60
2y-60=180-y
2y+y=180+60
3y=240
y=240/3
y=80 ans.
x=180-80
x=100 ans.
Proof:
100=2*80-60
100=160-60
100=100